# STaRK: Benchmarking LLM Retrieval on Textual and Relational Knowledge Bases

### v2

@@ 2.5 Synthesized Data Distribution Analysis and Human Evaluation
  • Query Diversity. A diverse set of queries pose challenges for broader applicability to meet varying user demands. Specifically, we measure query diversity using Shannon Entropy, which quantifies the uncertainty in the word distribution across all queries, and the Type-Token Ratio (TTR), which calculates the proportion of unique words to the total number of words. Higher values of Shannon Entropy and TTR indicate greater lexical diversity, reflecting a wider range of topics and linguistic expressions in the query set. As shown in Table 5, we observe high values of Shannon Entropy on all of the datasets. For reference, we compute the metrics for the wikipedia page of Barack Obama1footnotetext: 1https://en.wikipedia.org/wiki/Barack_Obama. While TTR is more sensitive to the document length, the TTR values consistently demonstrate a steady presence of unique words relative to the large length in our query document.
  • Ratio of Relational vs.Textual Information. A key feature of our benchmark dataset is the composition of textual and relational information. Therefore, it is crucial to understand the proportionality between these two types of information in the final queries. We calculate the ratio of relational to textual requirements for each query by comparing the word count of relational information to the word count of extracted textual properties, both of which are meta data in the constructing process. We then average these ratios across each dataset.
    Intuitively, the ratio serves as an approximation of the relative distribution based on the length of requirements, and does not directly reflect the importance of each type of information in determining the final answers. As shown in Figure 5, the ratios vary across different datasets, highlighting a differing emphasis on textual versus relational information. This distribution variation presents additional challenges for retrieval systems, requiring them to adapt to the dataset characteristic and specific balance of information.
Table 5: Positive/Non-negative rates (%) from human evaluation.
Naturalness Diversity Practicality
STaRK-Amazon 73.6 / 89.5 68.4 / 89.5 89.5 / 94.7
STaRK-MAG 94.7 / 100 73.7 / 84.2 68.4 / 84.2
STaRK-Prime 67.8 / 92.8 71.4 / 82.1 71.4 / 89.3
Average 78.7 / 94.1 71.0 / 85.3 76.4 / 89.4
@@
Human evaluation. Moreover, we qualitatively assess sampled queries from our benchmark to determine if they are natural-sounding (resembling natural conversation), diverse (covering a wide range of question structures and complexity levels), and practical (relevant and useful in real-life situations), with participation from 63 individuals. The evaluation results are converted from a 5-point Likert-like scale to a positive/tie/negative scale for reporting. We report the positive and non-negative rates in Table 5. On average across the three datasets, 94.1%, 85.3%, and 89.4% of participants rated neutral or above in terms of naturalness, diversity, and practicality, respectively. The results justify the quality of our benchmark and highlight its potential for diverse and realistic retrieval tasks.
@@ 2.6 Benchmark Construction: Human-Generated Queries
To enhance the practical relevance of our benchmark dataset, we engaged 31 participants, 22 of whom are native English speakers, to generate a total of 274 queries across three SKBs. The participants were given a list of entity IDs, randomly selected from the entire entity set. They followed detailed instructions (see Appendix D) using our interactive platform to explore the SKB data. Subsequently, we manually filtered the ground truth answers, ensuring their accuracy through human verification. We then analyze the commonality and difference between synthesized and human-generated queries.
@@
Complex Context Help me. I am trying to diagnose a patient with persistent joint pain, and I suspect a condition where the bone is dying due to compromised blood supply, often linked to factors like steroid use, alcohol abuse, or underlying diseases - what’s the name of this sneaky bone-killing culprit? I’m experiencing joint pain accompanied by swelling, fever… I’m concerned about medications aggravating my fuzzy eyesight and potential blood clotting complications. Could you recommend treatments effective for my symptoms while minimizing these specific side effects?
Commonality. We present the statistics of the human-generated datasets in Table 3. The synthesized and human-generated queries exhibit a similar level of ambiguity, resulting in a comparable number of answers. Moreover, we find that most styles and formats of human-generated queries are represented in our synthesized query dataset. In Table 6, we demonstrate three pairs of human-generated and synthesized queries that display similarities such as short and straightforward product queries, inquiries targeting specific authors and fields, and complex contextual queries in STaRK-Prime.
Difference. We find that human-generated queries often exhibit more unique expressions compared to synthesized queries, such as "Give me a fat cross and road tire that works with my Diamondback bicycle tube" and "this sneaky bone-killing culprit" (see the last row of Table 6). This discovery suggests a future direction for our benchmark to incorporate modern and dynamic language nuances.
@@ 3 Experiments
In this section, we aim to explore the following questions:
  • Q1: How well do current retrieval approaches perform on our benchmark datasets?
  • Q2: What challenges are revealed by the experiments, and more importantly, what are the potential future directions for improvement?
3.1 Retrieval Models
Vector Similarity Search (VSS). VSS embeds both the query and the concatenated textual and relational information of each candidate entity. Then, the similarity score is computed based on cosine similarity between the query and candidate embeddings. We use the text-embedding-ada-002 model from OpenAI for generating the embeddings.
Multi-Vector Similarity Search (Multi-VSS). Multi-VSS represents candidate entities with multiple vectors, capturing detailed features for complex retrieval tasks. Here, we chunk and separately embed the textual and relational information of each candidate entity. The final score is aggregated from the similarities between the chunks and the query.
Dense Retriever. Dense Retriever finetunes a query encoder and a document encoder separately using the query answer pairs from the training dataset. We optimize the encoder using contrastive loss. Specifically, the positive pairs are constructed from the training questions and their ground truth answers, while we construct 20 hard negative answers for each query from the top false positive predictions from VSS. We use roberta-base as the base model and finetune the encoder over the entire training split for each dataset.
QAGNN [49]. QAGNN constructs a graph where nodes include entities found in the question or answer choices, incorporating their neighboring nodes as intermediate context. It extracts a subgraph from a larger knowledge graph, enabling a comprehensive understanding by leveraging the relational information from the knowledge graph. The approach integrates this with semantic embeddings from a language model, jointly modeling both relational and semantic information to enhance the question-answering modeling.
VSS + LLM Reranker [9, 55]. This method improves the precision of the top-$v$ results from VSS by reranking them using language models, taking advantage of their advanced language understanding capabilities. For this purpose, we employ two different language models: GPT-4-turbo (gpt-4-1106-preview) and Claude3 (claude-3-opus). We set $v=20$ for synthesize queries and $v=10$ for human-generated queries to balance between precision and computational efficiency. Specifically, we construct a prompt that instructs the language models to assign a score between 0 and 1 to a node, based on its combined textual and relational information, with certain criteria provided for rating the node. Due to the high retrieval cost, we randomly sample 10% of the testing queries on each dataset for evaluation.
3.2 Evaluation Metrics
We use the following metrics for a holistically evaluating the model performance.
@@ 3.2 Evaluation Metrics
Hit@$k$. The Hit@$k$ metric assesses whether the correct item is among the top-$k$ results from the model. We used $k=1$ and $k=5$ for evaluation. At $k=1$, it evaluates the accuracy of the top recommendation; at $k=5$, it examines the model’s precision in a wider recommendation set. It’s a straightforward yet effective way to measure the immediate relevance of answers.
@@
Recall@$k$. Recall@$k$ measures the proportion of relevant items in the top-$k$ results. For synthesized queries, $k=20$ is used, as the answer length of all of the queries in our benchmarks are equal or smaller then $20$. For human-generated queries, Recall@20 is not reported as the number of responses in the human-generated subset may exceed 20. This metric offers insight into the model’s ability to identify all relevant items, particularly in scenarios where missing any could be critical.
@@
Mean Reciprocal Rank (MRR). MRR is a statistic for evaluating the average effectiveness of a predictive model. It calculates the reciprocal of the rank at which the first relevant item appears in the list of predictions. MRR is particularly useful for understanding the model’s performance at presenting the correct item at an early stage, without the need for a predefined $k$. This metric emphasizes the importance of the rank of the first correct answer, which is crucial in many practical applications where the first correct answer is often the most impactful.
3.3 Analysis
Table 7: Testing results on STaRK-Syn(thesized). (*) indicates partial evaluation (10% of the testing queries) due to the high latency and cost of the methods.
Dense Retriever QAGNN VSS Multi-VSS VSS+Claude3* VSS+GPT4*
(roberta) (roberta) (ada-002) (ada-002) Reranker Reranker
STaRK-Amazon Hit@1 0.1529 0.2656 0.3916 0.4007 0.4549 0.4479
Hit@5 0.4793 0.5001 0.6273 0.6498 0.7113 0.7117
Recall@20 0.4449 0.5205 0.5329 0.5512 0.5377 0.5535
MRR 0.3020 0.3775 0.5035 0.5155 0.5591 0.5569
STaRK-MAG Hit@1 0.1051 0.1288 0.2908 0.2592 0.3654 0.4090
Hit@5 0.3523 0.3901 0.4961 0.5043 0.5317 0.5818
Recall@20 0.4211 0.4697 0.4836 0.5080 0.4836 0.4860
MRR 0.2134 0.2912 0.3862 0.3694 0.4415 0.4900
STaRK-Prime Hit@1 0.0446 0.0885 0.1263 0.1510 0.1779 0.1828
Hit@5 0.2185 0.2135 0.3149 0.3356 0.3690 0.3728
Recall@20 0.3013 0.2963 0.3600 0.3805 0.3557 0.3405
MRR 0.1238 0.1473 0.2141 0.2349 0.2627 0.2655
@@ 3.3 Analysis
VSS Multi-VSS VSS+Claude3 VSS+GPT4
(ada-002) (ada-002) Reranker Reranker
STaRK-Amazon Hit@1 0.3950 0.4691 0.5602 0.5432
Hit@5 0.6419 0.7284 0.7284 0.7284
MRR 0.5265 0.5874 0.6439 0.6274
STaRK-MAG Hit@1 0.2857 0.2381 0.3810 0.3690
Hit@5 0.4167 0.4167 0.4643 0.4523
MRR 0.3581 0.3143 0.4243 0.4031
STaRK-Prime Hit@1 0.2120 0.2567 0.2844 0.2844
Hit@5 0.4037 0.4037 0.4771 0.4863
MRR 0.2984 0.3377 0.3623 0.3617
Results on Synthesized Queries. Table 7 shows the experiment results on the synthesized testing datasets. Firstly, we notice a significant performance gap between models trained or fine-tuned on our training datasets, such as Dense Retriever and QAGNN, and models like VSS and Multi-VSS that use embeddings generated from text-embedding-ada-002. For the dense retriever, training the encoders proves challenging when handling both textual and relational information as text, leading to training challenges. For QAGNN, it training challenges due to its high computational demands during training. Finally, the performance of VSS and Multi-VSS demonstrates the superior pre-trained embeddings obtained from a large corpus.
Further, the Multi-VSS strategy, which breaks down the documents into chunks and embeds them into embeddings separately to capture more fine-grained context, generally outperforms VSS. However, it falls short in the Hit@1 metric on the STaRK-MAG dataset, likely due to this dataset’s property for understanding more extended contexts or potential interaction or reasoning between different chunks.
Significantly, the methods combining VSS with an LLM reranker consistently outperform other approaches across most metrics. Notably, reranking the top 20 predictions with an LLM leads to significant improvements in Hit@$k$ and MRR metrics compared to using VSS alone. This improvement underscores the limitations of traditional embedding methods, which lack the advanced reasoning capabilities provided by LLMs.
However, even for VSS+GPT4 Reranker, we observe insufficient performances, e.g., Hit@1 is only around 18% on STaRK-Prime and 41% on STaRK-MAG , meaning that for a significant portion of queries, the top-ranked answer is not the correct one. Moreover, Recall@20 metric values are lower than 60% across all of the datasets, which indicates the ranking results may not be comprehensive, potentially missing relevant answers in specific applications. Also, we notice that the MRR scores generally follow a similar trend as the Hit@1 metric. However, the MRR values remain relatively low, particularly for STaRK-Prime, indicating large space for further optimization in the ranking process.
Results on human-generated dataset. We further evaluate the well-performing methods on human-generated queries. Table 8 presents the testing results on the human-generated dataset. Reranker methods continue to outperform other methods, confirming their superior reasoning and context understanding capabilities. Compared to synthesized datasets, the performance on human-generated queries shows highly correlated trends but generally higher metrics. This suggests that the synthesized datasets could be more difficult than the human-generated ones, offering challenging tasks and shedding light on interesting aspects of complex queries.
@@
Dense Retriever QAGNN VSS Multi-VSS VSS+Claude VSS+GPT4
@@
Retrieval latency. Another aspect vital for the practical application of retrieval systems is their latency. As shown in Table 9, we test the latency of the retrieval methods on a single NVIDIA A100-SXM4-80GB. Specifically, the Dense Retriever and QAGNN models exhibit lower average latency, making them more suitable for time-sensitive applications. In contrast, the VSS and Multi-VSS models show moderate latency, but when combined with LLM rerankers, such as Claude and GPT-4, the latency significantly increases. Therefore, in the context of semi-structured retrieval tasks, the balance between enhanced accuracy and increased latency is crucial. Effective retrieval methods should consider such trade-offs, particularly when addressing complex queries involving reasoning over both textual and relational information.
Figure 6: A case study on STaRK-MAG, where VSS mistakenly ranks non-ground truth papers C and D higher due to the repeated key words in the relational information "cites paper." After reranking with Claude, it correctly prioritizes the ground truth papers A and B. This correction is attributed to the more accurate reasoning and analysis of the combined textual and relational information.
Case study. Finally, we highlight the importance of reasoning ability in achieving good retrieval results on our benchmark datasets. In Figure 6, we conduct a case study to compare the performance of VSS and VSS+Claude3 Reranker. Specifically, we examine a user query that requests papers from an institution on a particular topic. In this scenario, VSS fails to adequately address the relational requirement due to directly embedding the entire documents without detailed analysis. As a result, papers containing frequently repeated keywords, such as “nonlinear modeling” and “piezoelectric elements”, are mistakenly assigned high scores. However, the result improves significantly after the reranking with an LLM, where the LLM is equipped to reason the relationship between the query and the information contained in each paper, making the scores more accurately reflect the papers’ relevance to the query. This case study underscores the need for reasoning ability in grasping the complexities of queries. As embedding methods may capture some aspects of the query, they often fail to capture the detailed interaction between textual and relational information. Thus, integrating LLMs with advanced reasoning into the retrieval process can be crucial for the retrieval accuracy.

### v3

@@ 2.5 Synthesized Data Distribution Analysis and Human Evaluation
  • Query Diversity. A diverse set of queries poses challenges for broader applicability to meet varying user demands. We measure query diversity using Shannon Entropy for word distribution and Type-Token Ratio (TTR) for unique words. Higher values indicate greater lexical diversity. Table 5 shows high Shannon Entropy and steady TTR across all datasets. For reference, we compute these metrics for the Wikipedia page of Barack Obamahttps://en.wikipedia.org/wiki/Barack_Obama.
  • Proportionality of Relational vs.Textual Information. Our benchmark queries feature the composition of textual and relational information. To understand the distribution of information types, we calculate the average ratio of relational to textual requirements by word count in the queries across each dataset. Note that the ratios do not directly reflect their importance in determining final answers. Figure 5 shows varying ratios, which highlights different emphases on textual versus relational requirements and challenges retrieval systems to adapt to different distributions.
@@
Human evaluation. We qualitatively assess sampled queries from our benchmark for naturalness (resembling natural conversation), diversity (covering various question structures and complexities), and practicality (relevance to real-world situations) with 63 participants. Evaluation results, converted from a 5-point Likert-like scale to a positive/tie/negative scale, show positive and non-negative rates in Table 10 (Appendix D.1). On average, 94.1%, 85.3%, and 89.4% of participants rated the queries neutral or above in naturalness, diversity, and practicality, respectively. These results validate the quality of our benchmark and its potential for diverse and realistic retrieval tasks.
@@ 2.6 Benchmark Construction: Human-Generated Queries
To enhance our benchmark’s practical relevance, we engaged 31 participants (22 native English speakers) to generate 263 queries across three SKBs following the detailed instructions (Appendix C) along with our interactive platform. We manually verified and filtered the ground truth answers to ensure the answer correctness. Table 3 shows the statistics of the human-generated datasets. Finally, we analyzed the commonalities and differences between synthesized and human-generated queries.
@@
Complex Context Help me. I am trying to diagnose a patient with persistent joint pain, and I suspect a condition where the bone is dying due to compromised blood supply, often linked to factors like steroid use, … what’s the name of this sneaky bone-killing culprit? I’m experiencing joint pain accompanied by swelling… I’m concerned about medications aggravating my fuzzy eyesight and potential blood clotting complications. Could you recommend treatments while minimizing these side effects?
Commonality. The number of answers of synthesized and human-generated queries are comparable, indicating a similar level of query ambiguity. Moreover, we observe that most styles of human-generated queries are covered in the synthesized dataset. For example, Table 5 highlights their similarities in short product queries, specific author/field inquiries, and complex contextual queries.
Difference. We find that human-generated queries often exhibit more unique expressions compared to synthesized ones, such as "Give me a fat cross and road tire that works with my Diamondback bicycle tube" and "this sneaky bone-killing culprit". This discovery suggests a future direction for our benchmark to incorporate modern and dynamic language nuances.
@@ 3 Experiments
3.1 Baseline Retrieval Models and Evaluation Metrics
We extensively evaluate five classes of retrieval models described below.
  • Sparse Retriever: BM25 [40] is a traditional yet powerful sparse retrieval method based on term frequency-inverse document frequency (TF-IDF). It computes relevance scores by considering the frequency of query terms in documents, adjusted for term rarity and document length.
  • Small Dense Retrievers: DPR [27], ANCE [53], and QAGNN [57]. These compact models generate dense embeddings for both queries and documents, computing retrieval scores based on embedding similarities. They serve as baselines for comparison with LLM-based dense retrievers.
  • LLM-based Dense Retrievers: text-embedding-ada-002 (abbrev. ada-002) [37], voyage-large-2-instruct (abbrev. voyage-l2-instruct) [2], LLM2Vec-Meta-Llama-3-8B-Instruct-mntp (abbrev. LLM2Vec) [4], and GritLM-7b [34]. These models leverage LLMs to generate dense embeddings that are more contextually expressive.
  • Multivector Retrievers: multi-ada-002 [37] and ColBERTv2 [42]. Beyond ada-002 which represents a document as an embedding, multi-ada-002 splits each document into overlapping chunks and embeds them using the same encoder as the query. Similarity scores between the query and chunks are aggregated using the average of the top-3 similarities, which we found to perform best. ColBERTv2 represents each document as multiple token-level embeddings for fine-grained matching, capturing richer semantic information.
  • LLM Rerankers: Claude3 and GPT-4 rerankers [12, 63]. These models improve the precision of top-$k$ ada-002 results by reranking them using large language models. We employ GPT-4-turbo (gpt-4-1106-preview) and Claude3 (claude-3-opus), setting $k=20$ for synthesized queries and $k=10$ for human-generated queries. Given a query, the LLMs assign a satisfaction score from 0 to 1 to each candidate entity based on textual and relational information. Due to high computational costs, we evaluate these rerankers on a random 10% sample of test queries.
The performance of these models are measured using standard retrieval metrics below.
@@ 3.1 Baseline Retrieval Models and Evaluation Metrics
  • Hit@$k$ assesses whether the correct item is among the top-$k$ results from the model. We used $k=1$ and $k=5$ for evaluation. At $k=1$, it evaluates the accuracy of the top recommendation; at $k=5$, it examines the model’s precision in a wider recommendation set.
@@
  • Recall@$k$ measures the proportion of relevant items in the top-$k$ results. For synthesized queries, $k=20$ is used, as the answer length of all of the queries in our benchmarks are equal or smaller then $20$. This metric offers insight into the model’s ability to identify all relevant items, particularly in scenarios where missing any could be critical.
@@
  • Mean Reciprocal Rank (MRR) is a statistic for evaluating the average effectiveness of a predictive model. It calculates the reciprocal of the rank at which the first relevant item appears in the list of predictions. This metric emphasizes the importance of the rank of the first correct answer, which is crucial in many practical applications where the first correct answer is often the most impactful.
3.2 Results and Analysis
Table 6: Testing results on STaRK-Syn(thesized).
STaRK-Amazon STaRK-MAG STaRK-Prime
Hit@1 Hit@5 R@20 MRR Hit@1 Hit@5 R@20 MRR Hit@1 Hit@5 R@20 MRR
Full Testing Dataset
BM25 44.94 67.42 53.77 55.30 25.85 45.25 45.69 34.91 12.75 27.92 31.25 19.84
DPR (roberta) 15.29 47.93 44.49 30.20 10.51 35.23 42.11 21.34 4.46 21.85 30.13 12.38
ANCE (roberta) 30.96 51.06 41.95 40.66 21.96 36.50 35.32 29.14 6.53 15.67 16.52 11.05
QAGNN (roberta) 26.56 50.01 52.05 37.75 12.88 39.01 46.97 29.12 8.85 21.35 29.63 14.73
ada-002 39.16 62.73 53.29 50.35 29.08 49.61 48.36 38.62 12.63 31.49 36.00 21.41
voyage-l2-instruct 40.93 64.37 54.28 51.60 30.06 50.58 50.49 39.66 10.85 30.23 37.83 19.99
LLM2Vec 21.74 41.65 33.22 31.47 18.01 34.85 35.46 26.10 10.10 22.49 26.34 16.12
GritLM-7b 42.08 66.87 56.52 53.46 37.90 56.74 46.40 47.25 15.57 33.42 39.09 24.11
multi-ada-002 40.07 64.98 55.12 51.55 25.92 50.43 50.80 36.94 15.10 33.56 38.05 23.49
ColBERTv2 46.10 66.02 53.44 55.51 31.18 46.42 43.94 38.39 11.75 23.85 25.04 17.39
Random 10% Sample
BM25 42.68 67.07 54.48 54.02 27.81 45.48 44.59 35.97 13.93 31.07 32.84 21.68
DPR (roberta) 16.46 50.00 42.15 30.20 11.65 36.84 42.30 21.82 5.00 23.57 30.50 13.50
ANCE (roberta) 30.09 49.27 41.91 39.30 22.89 37.26 44.16 30.00 6.78 16.15 17.07 11.42
QAGNN (roberta) 25.00 48.17 51.65 36.87 12.03 37.97 47.98 28.70 7.14 17.14 32.95 16.27
ada-002 39.02 64.02 49.30 50.32 28.20 52.63 49.25 38.55 15.36 31.07 37.88 23.50
voyage-l2-instruct 43.29 67.68 56.04 54.20 34.59 50.75 50.75 42.90 12.14 31.42 37.34 21.23
LLM2Vec 18.90 37.80 34.73 28.76 19.17 33.46 29.85 26.06 9.29 20.7 25.54 15.00
GritLM-7b 43.29 71.34 56.14 55.07 38.35 58.64 46.38 48.25 16.79 34.29 41.11 24.99
multi-ada-002 40.85 62.80 52.47 51.54 25.56 50.37 53.03 36.82 15.36 32.86 40.99 23.70
ColBERTv2 44.31 65.24 51.00 55.07 31.58 47.36 45.72 38.98 15.00 26.07 27.78 19.98
Claude3 Reranker 45.49 71.13 53.77 55.91 36.54 53.17 48.36 44.15 17.79 36.90 35.57 26.27
GPT4 Reranker 44.79 71.17 55.35 55.69 40.90 58.18 48.60 49.00 18.28 37.28 34.05 26.55
@@ 3.2 Results and Analysis
STaRK-Amazon STaRK-MAG STaRK-Prime
Method Hit@1 Hit@5 R@20 MRR Hit@1 Hit@5 R@20 MRR Hit@1 Hit@5 R@20 MRR
BM25 27.16 51.85 29.23 18.79 32.14 41.67 32.46 37.42 22.45 41.84 42.32 30.37
DPR (roberta) 16.05 39.51 15.23 27.21 4.72 9.52 25.00 7.90 2.04 9.18 10.69 7.05
ANCE (roberta) 25.93 54.32 23.69 37.12 25.00 30.95 27.24 27.98 7.14 13.27 11.72 10.07
QAGNN (roberta) 22.22 49.38 21.54 31.33 20.24 26.19 28.76 25.53 6.12 13.27 17.62 9.39
ada-002 39.50 64.19 35.46 52.65 28.57 41.67 35.95 35.81 17.35 34.69 41.09 26.35
voyage-l2-instruct 35.80 62.96 33.01 47.84 22.62 36.90 32.44 29.68 16.33 32.65 39.01 24.33
LLM2Vec 29.63 46.91 21.21 38.61 16.67 28.57 21.74 21.59 9.18 21.43 26.77 15.24
GritLM-7b 40.74 71.60 36.30 53.21 34.52 44.04 34.57 38.72 25.51 41.84 48.10 34.28
multi-ada-002 46.91 72.84 40.22 58.74 23.81 41.67 39.85 31.43 24.49 39.80 47.21 32.98
ColBERTv2 33.33 55.56 29.03 43.77 33.33 36.90 30.50 35.97 15.31 26.53 25.56 19.67
Claude3 Reranker 53.09 74.07 35.46 62.11 38.10 45.24 35.95 42.00 28.57 46.94 41.61 36.32
GPT4 Reranker 50.62 75.31 35.46 61.06 36.90 46.43 35.95 40.65 28.57 44.90 41.61 34.82
Results on synthesized queries. Table 6 presents the results on both the full synthesized test sets and random 10% samples from these sets. In both cases, BM25, despite its simplicity, proves to be a strong baseline, outperforming the dense retrieval models such as ANCE. We observe that finetuned DPR and QAGNN, exhibit insufficient performance. This underperformance is likely due to their relatively small model sizes and the risk of overfitting during training. These issues present challenges in effectively training the models on SKBs, where the entity documents can be hard to differentiate without capturing detailed information.
Among the larger models, ada-002 benefits from superior pretrained embeddings and significantly outperforms LLM2Vec by a large margin. GritLM-7b delivers excellent performance, surpassing the ada-002 model overall. In contrast, LLM2Vec underperforms due to its limited context length, which is insufficient for encoding the lengthy documents in the SKBs. For multivector retrievers, we found that multi-ada-002 generally outperforms ada-002, indicating that using multiple vectors per document enhances retrieval effectiveness. Similarly, fine-grained representation allows ColBERTv2 to capture subtle semantic nuances between queries and documents, leading to largely improved retrieval accuracy.
However, both GritLM-7b and ColBERTv2 generally underperform compared to the rerankers on the random split, especially in terms of Hit@k metrics. This suggests that while these dense retriever models effectively capture semantic information, they may not fully grasp the nuanced relevance judgments required for top-tier retrieval performance. The rerankers, utilizing powerful LLMs like GPT-4 (gpt-4-1106-preview) and Claude3 (claude-3-opus), excel by re-evaluating the top candidates and assigning satisfaction scores based on a deeper understanding of the query and document content. This process allows them to better discern subtle contextual cues and relational information that dense retrievers might overlook. Consequently, LLM rerankers enhance retrieval precision at the top ranks.
Finally, regardless of the higher computational costs of the rerankers, their performance remains suboptimal. For instance, the Hit@1 scores for the GPT-4 reranker are only about 18% on STaRK-Primeand 41% on STaRK-MAG, indicating that the top-ranked answers are frequently incorrect. Similarly, the Recall@20 metrics are below 60% across all datasets, with the GPT-4 reranker achieving Recall@20 scores of 55% on STaRK-Amazon, 49% on STaRK-MAG, and 34% on STaRK-Prime. This suggests that the ranking results miss a significant portion of relevant answers. The MRR scores are also relatively low, especially for STaRK-Prime, where the GPT-4 reranker attains an MRR of only around 27%.
The insufficient performances may be attributed to the complexity and diversity of queries in SKBs, where nuanced understanding and detailed contextual information are crucial. These findings highlight significant room for improvement in the ranking process.
Results on human-generated dataset. Table 7 presents the testing results on the human-generated datasets. For example, the rerankers consistently outperform others, showing their reasoning and context understanding ability. Compared to the synthesized datasets, the performance on human-generated queries is generally higher for most models, but the overall trends remain consistent. This indicates that synthesized datasets may be more challenging, highlighting the complexity of the tasks on our synthesized queries.
Another interesting observation is that the performance of the rerankers is particularly strong on human-generated queries, which may contain more nuanced language and diverse expressions. This suggests that rerankers excel in interpreting and leveraging the richness of human language to improve retrieval accuracy.
@@
DPR QAGNN ada-002 multi-ada-002 Claude3 Reranker GPT4 Reranker
@@
Figure 6: A case study on STaRK-MAG shows that ada-002 overranks non-ground truth papers $C$ and $D$ due to repeated keywords in the relational information “cites paper”. After reranking with Claude3, it correctly prioritizes ground truth papers $A$ and $B$ with accurate reasoning and analysis.
Retrieval latency. Latency is crucial for practical retrieval systems, as users expect quick responses. As shown in Table 8, we evaluated the latency of various models using a single NVIDIA A100-SXM4-80GB GPU. We observed that the DPR and QAGNN models exhibit lower average latency, making them suitable for time-sensitive applications. In contrast, the ada-002 and multi-ada-002 models have moderate latency due to multiple API calls. However, when combined with LLM rerankers, the latency increases significantly due to the computational demands of these large models. Therefore, it is important to balance accuracy and latency, especially for complex queries that require advanced reasoning capabilities.
Case study. To highlight the importance of reasoning ability for achieving good performance on our benchmark, we present a case study in Figure 6, comparing the ada-002 model with the Claude3 Reranker. In this example, the query requests papers from a specific institution on a particular topic. The ada-002 model fails to address the relational aspect of the query because it embeds entire documents without detailed analysis. This leads to high relevance scores for irrelevant papers that frequently mention keywords like "nonlinear modeling" and "piezoelectric elements" but do not satisfy the relational requirement. In contrast, the LLM reranker significantly improves the results by reasoning about the relationship between the query and each paper, resulting in scores that more accurately reflect relevance. This underscores the need for reasoning ability to grasp query complexities.